OpenStudy (anonymous):

after 100 days a particular radioactive substance decays to 30% of its original amount. find the number of days it will take for this substance to decay to 60% of its original amount

OpenStudy (anonymous):

Are you given the original amount?

OpenStudy (anonymous):

no. that's the question

OpenStudy (anonymous):

I'm assuming the original amount is 100%

OpenStudy (amistre64):

good assumption

OpenStudy (amistre64):

30 = 100 e^(100r) solve for r

OpenStudy (anonymous):

doesn't matter sovle \[(.3)^{\frac{t}{100}}=.6\] for t

OpenStudy (anonymous):

or use amsitre method, find r, and then go back and solve for t either way

OpenStudy (anonymous):

\[\frac{t}{100}=\frac{\ln(.6)}{\ln(.3)}\] \[t=\frac{100\ln(.6)}{\ln(.3)}\] then a calculator

OpenStudy (anonymous):

thank you guys!

OpenStudy (anonymous):

42.4

OpenStudy (anonymous):

since u can't have 42.4 days you round it off to 42 days.

OpenStudy (anonymous):

if i use amistre 64's method...how would i go back and solve for t?

OpenStudy (rogue):

Amistre's method: \[0.3 = e^{100r} \rightarrow r = \frac {1}{100} \ln 0.3\]\[60 = 100e^{\frac {t}{100} \ln 0.3} \rightarrow 0.6 = e^{\frac {t}{100} \ln 0.3} \rightarrow t = \frac {100 \ln 0.6}{\ln 0.3} \approx 42.428 days\]

OpenStudy (anonymous):

thank u

OpenStudy (rogue):

welcome =)